OpenStudy (anonymous):

If T is the number of teaspoons, then 1 _ 3 T can be used to find the number of tablespoons. Find the tablespoons of spice in each recipe. 5. 3 teaspoons 6. 2 teaspoons 7. 4 teaspoons 8. 7 teaspoons can anyone help?

OpenStudy (noelgreco):

What does 1_3 mean? If it's 1/3, then it's wrong: there are three teaspoons in a Tablespoon.

OpenStudy (noelgreco):

You simply have to multiply the number of teaspoons by \[\frac{ 1 }{ 3 }\]

OpenStudy (noelgreco):

Multiplying fractions is simple; \[\frac{ 1 }{ 3 }\times \frac{ 7 }{ 1 }=\frac{ 7 }{ 3 }\] Just multiply the numerators and the denominators